OpenStudy (anonymous):

Solve the system x + y = 6 and x - y = 2. What is the value of x? A. 6 B. 2 C. 4 D. 0

OpenStudy (swag):

A

OpenStudy (vishal_kothari):

http://openstudy.com/code-of-conduct

OpenStudy (vishal_kothari):

x + y = 6 x - y = 9 Add the two equations: 2x = 15 x = 15/2 Substitute x = 15/2 in any of the two equations: x - y = 9 15/2 - y = 9 y = - 3/2...